OpenStudy (anonymous):

A school system is reducing the amount of dumpster loads of trash removed each week. In week 5, there were 40 dumpster loads of waste removed. In week 10, there were 30 dumpster loads removed. Assume that the reduction in the amount of waste each week is linear. Write an equation in function form to show the amount of trash removed each week.

OpenStudy (anonymous):

Giving medals

OpenStudy (anonymous):

f(x) = - 2x + 40 f(x) = 2x + 40 f(x) = - 2x + 50 f(x) = 2x + 50

OpenStudy (anonymous):

really need help!!

OpenStudy (damonte212):

y=-2x+50 is the answer

OpenStudy (anonymous):

So would it be f(x) = -2x + 50 ???

OpenStudy (anonymous):

@damonte212

OpenStudy (anonymous):

So would it be f(x) = -2x + 50 ???

OpenStudy (damonte212):

exactly

OpenStudy (anonymous):

Thank you, you get a medal!!!!

OpenStudy (damonte212):

thanks and jus tag me I'm here

OpenStudy (anonymous):

okay, thanks

OpenStudy (anonymous):

Can you help with one more please?

OpenStudy (damonte212):

jus tag me